# Learning Protein-Ligand Binding in Hyperbolic Space

### v1

@@ 5.2 Quantitative Results
Glide-SP Friesner et al. (2004) 0.7670 0.4070 16.18 0.5315 0.4000 3.41
Surflex Spitzer & Jain (2012) 0.7426 0.2387 13.35 0.5147 2.50
DeepDTA Öztürk et al. (2018) 0.5836 0.0513 2.28 0.5627 0.0253 1.47
Gnina McNutt et al. (2021a) 0.7817 0.2994 17.73 0.6093 0.0540 4.63
BigBind Brocidiacono et al. (2022) 0.5014 0.0240 1.18 0.6278 0.0502 3.79
RTMScore Shen et al. (2022) 0.7529 0.4341 27.10 0.5247 0.0388 2.94
Tankbind Lu et al. (2022) 0.7509 0.3300 13.00 0.5970 0.0389 2.90
DrugCLIP Gao et al. (2023a) 0.8093 0.5052 31.89 0.5717 0.0623 5.51
GenScore Shen et al. (2023) 0.8160 0.4726 28.53 0.5957 0.0654 5.14
Planet Zhang et al. (2024) 0.7160 8.83 0.5731 3.87
EquiScore Cao et al. (2024) 0.7760 0.4320 17.68 0.5678 0.0490 3.51
DrugHash Han et al. (2025b) 0.8373 0.5716 37.18 0.5458 0.0714 6.14
LigUnitypoc Feng et al. (2025) 0.8922 0.6526 42.63 0.5985 0.1133 6.47
@@
Graph CNN Torng & Altman (2019) 0.8860 44.41 29.75 19.41 10.74
DrugVQA Zheng et al. (2020) 0.9720 88.17 58.71 35.06 17.39
AttentionSiteDTI Yazdani-Jahromi et al. (2022) 0.9710 101.74 59.92 35.07 16.74
COSP Gao et al. (2023b) 0.9010 51.05 35.98 23.68 12.21
DrugCLIPZS Gao et al. (2023a) 0.8093 73.97 41.79 23.68 11.16
DrugCLIPFT Gao et al. (2023a) 0.9659 118.10 67.17 37.17 16.59
LigUnitypoc Feng et al. (2025) 0.8922 104.69 57.47 33.76 13.88
@@
Physics FEP+ Wang et al. (2015) 0.7811 0.7595 0.6960 0.6798
MM-GB/SA Genheden & Ryde (2015) 0.1489 0.2011 0.1299 0.1299
DL PBCNet Yu et al. (2023) 0.3939 0.3799 0.4058 0.4075
EHIGN Yang et al. (2024) 0.5787 0.5814 0.4246 0.3830
GET Kong et al. (2024) 0.4034 0.3753 0.4203 0.4214
BindNet Feng et al. (2024) 0.5481 0.5368 0.4037 0.3477
Boltz-2 Passaro et al. (2025) 0.5231 0.5285 0.4298 0.4013
LigUnitypoc (ensemble) Feng et al. (2025) 0.6454 0.6460 0.5997 0.5554
@@ 5.3 Ablation and Analysis of HypSeek
Pairwise Affinity Prediction. Figure 3 (A)-(B) demonstrate the behavior of Euclidean and hyperbolic models across varying ECFP4 Rogers & Hahn (2010) similarity. Both models perform similarly on dissimilar ligand pairs, but as the ligands become more structurally similar, Euclidean accuracy and correlation decrease significantly. In contrast, the hyperbolic model maintains strong performance, even in these highly similar pairs. This suggests that the richer geometry information in hyperbolic space, which better accommodates relationships between molecules, is more effective at capturing subtle affinity shifts typical of situations where structurally similar molecules exhibit significantly different biological activity. These differences are often compressed in Euclidean space, where the geometry may fail to distinguish between such subtle shifts.
Embedding Visualization. Ligand embeddings are first reduced via HoroPCA Chami et al. (2021) and visualized using CO-SNE Guo et al. (2022). Without hyperbolic constraints (Figure 3C), embeddings collapse near the origin with overlapping targets. With the full HypSeek objective (Figure 3D), clear target-wise clusters and radial affinity gradients emerge. This contrast illustrates how the cone–hierarchy constraints introduced by HypSeek structure the hyperbolic manifold, enabling a more effective representation of the complex relationships between ligands in hyperbolic space.
@@ B.1 Analysis of Cross-Target Activity–Cliff Pairs
Table 6 lists 21 ligand pairs whose ECFP Rogers & Hahn (2010) similarity is greater than 0.60 yet display large differences in experimental binding free energy Exp $\Delta G$  making them representative activity–cliff cases for evaluating our embedding space. For comparison, the Euclidean scores in the table are produced by the current state-of-the-art pocket–ligand model LigUnitypoc Feng et al. (2025), whereas the hyperbolic scores come from our method.
Directional Agreement with Experimental Affinity. Recall that a smaller (more negative) experimental $\Delta G$ indicates a stronger binder, whereas a larger model score indicates stronger binding. Hence, for every pair in Table 6 we expect the sign of $\Delta(\text{score})$ to be opposite to the sign of $\Delta(\text{Exp}\,\Delta G)$. This correspondence is clearly visible: whenever the experimental gap favours molecule B, the hyperbolic score is higher for B (positive $\Delta\text{Hyp}$), and vice-versa. Euclidean scores occasionally match the sign but the margin is often negligible. Several pairs show that even free-energy perturbation (FEP) Wang et al. (2015) predicts the wrong direction of the affinity change, yet the hyperbolic score still aligns with the experimental ordering.

### v2

@@ 5.2 Quantitative Results
Glide-SP friesner_glide_2004 0.7670 0.4070 16.18 0.5315 0.4000 3.41
Surflex spitzer2012surflex 0.7426 0.2387 13.35 0.5147 2.50
DeepDTA ozturk2018deepdta 0.5836 0.0513 2.28 0.5627 0.0253 1.47
Gnina mcnutt2021gnina 0.7817 0.2994 17.73 0.6093 0.0540 4.63
BigBind brocidiacono2022bigbind 0.5014 0.0240 1.18 0.6278 0.0502 3.79
RTMScore shen_rtmscore_2022 0.7529 0.4341 27.10 0.5247 0.0388 2.94
Tankbind lu_tankbind_2022 0.7509 0.3300 13.00 0.5970 0.0389 2.90
DrugCLIP gao2023drugclip 0.8093 0.5052 31.89 0.5717 0.0623 5.51
GenScore shen2023generalized 0.8160 0.4726 28.53 0.5957 0.0654 5.14
Planet zhang_planet_2024 0.7160 8.83 0.5731 3.87
EquiScore cao_equiscore_2024 0.7760 0.4320 17.68 0.5678 0.0490 3.51
DrugHash Han_Hong_Li_2025 0.8373 0.5716 37.18 0.5458 0.0714 6.14
LigUnitypoc feng2025foundation 0.8922 0.6526 42.63 0.5985 0.1133 6.47
@@
Graph CNN doi:10.1021/acs.jcim.9b00628 0.8860 44.41 29.75 19.41 10.74
DrugVQA Zheng2020 0.9720 88.17 58.71 35.06 17.39
AttentionSiteDTI 10.1093/bib/bbac272 0.9710 101.74 59.92 35.07 16.74
COSP Gao2023CoSP 0.9010 51.05 35.98 23.68 12.21
DrugCLIPZS gao2023drugclip 0.8093 73.97 41.79 23.68 11.16
DrugCLIPFT gao2023drugclip 0.9659 118.10 67.17 37.17 16.59
LigUnitypoc feng2025foundation 0.8922 104.69 57.47 33.76 13.88
@@
Physics FEP+ wang2015accurate 0.7811 0.7595 0.6960 0.6798
MM-GB/SA genheden2015mm 0.1489 0.2011 0.1299 0.1299
DL PBCNet yu2023computing 0.3939 0.3799 0.4058 0.4075
EHIGN yang2024interaction 0.5787 0.5814 0.4246 0.3830
GET pmlr-v235-kong24b 0.4034 0.3753 0.4203 0.4214
BindNet feng2024proteinligand 0.5481 0.5368 0.4037 0.3477
Boltz-2 Passaro2025.06.14.659707 0.5231 0.5285 0.4298 0.4013
LigUnitypoc (ensemble) feng2025foundation 0.6454 0.6460 0.5997 0.5554
@@ 5.3 Ablation and Analysis of HypSeek
Pairwise Affinity Prediction. Figure 3 (A)-(B) demonstrate the behavior of Euclidean and hyperbolic models across varying ECFP4 doi:10.1021/ci100050t similarity. Both models perform similarly on dissimilar ligand pairs, but as the ligands become more structurally similar, Euclidean accuracy and correlation decrease significantly. In contrast, the hyperbolic model maintains strong performance, even in these highly similar pairs. This suggests that the richer geometry information in hyperbolic space, which better accommodates relationships between molecules, is more effective at capturing subtle affinity shifts typical of situations where structurally similar molecules exhibit significantly different biological activity. These differences are often compressed in Euclidean space, where the geometry may fail to distinguish between such subtle shifts.
Embedding Visualization. Ligand embeddings are first reduced via HoroPCA pmlr-v139-chami21a and visualized using CO-SNE 9879850. Without hyperbolic constraints (Figure 3C), embeddings collapse near the origin with overlapping targets. With the full HypSeek objective (Figure 3D), clear target-wise clusters and radial affinity gradients emerge. This contrast illustrates how the cone–hierarchy constraints introduced by HypSeek structure the hyperbolic manifold, enabling a more effective representation of the complex relationships between ligands in hyperbolic space.
@@ 6 Conclusion
7 Acknowledgments
This work was supported by the National Key R&D Program of China No. 2025ZD1802501, the National Natural Science Foundation of China (NSFC) No. 62506193, and the Beijing Academy of Artificial Intelligence.
@@ B.1 Analysis of Cross-Target Activity–Cliff Pairs
Table 6 lists 21 ligand pairs whose ECFP doi:10.1021/ci100050t similarity is greater than 0.60 yet display large differences in experimental binding free energy Exp $\Delta G$  making them representative activity–cliff cases for evaluating our embedding space. For comparison, the Euclidean scores in the table are produced by the current state-of-the-art pocket–ligand model LigUnitypoc feng2025foundation, whereas the hyperbolic scores come from our method.
Directional Agreement with Experimental Affinity. Recall that a smaller (more negative) experimental $\Delta G$ indicates a stronger binder, whereas a larger model score indicates stronger binding. Hence, for every pair in Table 6 we expect the sign of $\Delta(\text{score})$ to be opposite to the sign of $\Delta(\text{Exp}\,\Delta G)$. This correspondence is clearly visible: whenever the experimental gap favours molecule B, the hyperbolic score is higher for B (positive $\Delta\text{Hyp}$), and vice-versa. Euclidean scores occasionally match the sign but the margin is often negligible. Several pairs show that even free-energy perturbation (FEP) wang2015accurate predicts the wrong direction of the affinity change, yet the hyperbolic score still aligns with the experimental ordering.